OpenStudy (anonymous):

Write an equation of the direct variation that includes the given point. (1, 4)

OpenStudy (anonymous):

What is the definition of direct variation?

OpenStudy (anonymous):

@TylerD @adrynicoleb

OpenStudy (anonymous):

Direct variation means that x and y are directly proportional to each other, hence, if you double x, y is doubled as well. Direct variation equations will have the form \[y = ax\]

OpenStudy (anonymous):

y=14?

OpenStudy (anonymous):

so y = 4x?

OpenStudy (tylerd):

possibly

OpenStudy (anonymous):

what about (-6, -9)?

OpenStudy (triciaal):

y = k x where k is a constant will not change using (1, 4) 4 = k *1 k = 4 therefore y = 4 *x

OpenStudy (anonymous):

what about (-6, -9)?

OpenStudy (tylerd):